# Proactive Agents for Multi-Turn Text-to-Image Generation Under Uncertainty

### v1

@@ 5.3 Analysis of quantitative results
Table 2: Perceived helpfulness of proposed features (of users) rated by 143 raters.
5.4 Human studies on generated images and dialogues
To verify the automatic evaluations of the agents, we performed human studies in which participants were asked to rate the generated images and dialogues along different axes. The detailed design of the studies can be found in Figure 20, Figure 21 and Figure 22.
Participants are asked to rate the images produced by the three proposed multi-turn agents and a single-turn T2I model against a Ground Truth image for which the original prompt was derived and the answers to the agents questions were derived. Approximately 550 image-dialog pairs per agent are rated using 3 human raters. The generated images were presented in a random order and were unlabeled and the human rater was tasked with ranking the images from best to worst. The results from the study are shown in Figure 4. We used 3 raters per set of images and therefore in cases where raters did not agree this has been noted via the No Agreement Column. The graph shows that the agentic systems are selected as the best generated image over the single-turn T2I in 80%+ of cases for both content and style.
Figure 4: Human Rating of the Generated Images. Ratings are based on Content Correctness and Style and Aesthetics. Each human rater is given the Ground Truth Image and Prompt to compare the Generated Image against.
To validate the generated dialog, human raters are asked to mark any issues a question contains that could pose a disturbance to the user. Approximately 8k questions per agent are rated. The results are shown in Figure 5, where we see that the agentic systems have issues with their questions in 14% or less cases. For the Ag2 and Ag3 the common complaint is that the question is too long while the most common issue for Ag1 is that the question does not gain any new information.
Human raters are also asked to rank the correspondence of each image to the agent-user dialog and original prompt. Approximately 1.5k image-dialog pairs are rated using 3 human raters. Results in Figure 5 show that for all of our agents, more than 96% of the 1.5k image-dialog pairs are rated as very close or fairly close with some differences. This high rating shows the viability of the T2I model employed by the agents, as well as the agents’ ability to combine the dialogue into a coherent prompt to feed the T2I model.
Figure 5: Human ratings for the dialogues. The left graph shows the rating of how well the final generated image corresponds to the original prompt and dialogue. The right graph shows the distribution of issues of questions per agent.
5.5 Human studies on the agent interface
To get real user feedback on the agent interface, we performed a human survey with the objective of understanding user frustrations and validating our solutions. We gathered data from 143 participants who all identified to be regular T2I users (at least once a month). Participants were presented with four hypothesized frustrations (prompt misinterpretation, many iterations, inconsistent generations, incorrect assumptions) and three potential mitigating features (clarifications, entity graph, relationship graph; more details in §H).
Table 4 in Appendix confirms the prevalence of hypothesized frustrations amongst users, with 83% experiencing occasional, frequent, or very frequent frustration due to prompt iterations, followed by 70% for misinterpretations, 71% for inconsistent generations, and 60% experiencing frustration due to incorrect assumptions. Most acutely 55% of participants reported frequent or very frequent frustration due to the prompt iteration frequency necessary. In Table 2, we report the mitigation features that are likely to help. Clarifications reported the highest likelihood to help current workflows (91% could / likely / very likely to be helpful), followed by entity graphs (88% could / likely / very likely to be helpful) and relationship graphs (86% could / likely / very likely to be helpful). Clarifications were expected to deliver value immediately / very soon by 58%.
Overall these suggest strong user desire for & likelihood for success of features that reduce iterations and mitigate misinterpretations in T2I generation. Full explanations of the hypothesized frustrations, mitigation and responses splits are in §H. All respondents were compensated for their time as per market rates, and were recruited by our vendor to ensure diversity across age, gender, and T2I usage in terms of models, frequency and purpose (work and non work).
@@ 6 Discussion and conclusion
This work introduces a design for agents that assist users in generating images through an interactive process of proactive question asking and belief graph refinement. By dynamically updating its understanding of the user’s intent, the agent facilitates a more collaborative and precise approach to image generation. Moreover, presenting the agent’s belief graph can be a generalizable method for AI transparancy, which is an important factor given the increasing complexity of modern AI models.
Modular design. Our agent prototypes are highly modular: the agents use frozen T2I models to generate images based on the prompts that the agent updated. Therefore when a better off-the-shelf T2I model becomes available, it can be directly plugged into the agents and the system will achieve better performance without any additional adaptation [^9] T2T scores in Table 1 ablates the T2I model and only performs similarity on the captions. Our agents have achieved a 92%+ T2T score, showing that their performance can be boosted by adopting better T2I models..
Personalized content. By asking clarification questions, our agents enable a more customizable and personalized content creation experience. Because different groups of people may perceive helpfulness and harmfulness of contents differently, learning more about the user through clarification questions before generation can potentially mitigate risks of generating contents that are offensive to each specific user, and increase likelihoods of producing helpful outputs.
Future work. Alternative to the modular design, one can explore generating images directly from belief graphs and fine-tuning LLM/VLMs on text/image trajectories that include asking questions. These may require a) collecting data such as gold-standard trajectories or annotations on the quality of trajectories of human-agent conversations and b) new approaches to fine-tune the model on multi-turn trajectories of images and text, which can potentially improve the performance of the agent.
@@ Acknowledgements
We would like to thank Jason Baldridge and Zoubin Ghahramani for insightful discussions on multi-turn T2I and belief states, Mahima Pushkarna for the help and consultation on user study. We would also like to thank Richard Song and Noah Fiedel for feedback on the paper.
@@ Appendix A Author contributions
Appendix B Novelty and contributions
In this section, we emphasize the novelty and contributions of this work.
  1. 1.
    System design of proactive T2I agents.
    • Novel human-agent interaction modalities: Prior to our work, human users typically interact with current T2I systems by giving additional instructions or refining the prompt. To the best of our knowledge, our work is the first to propose a proactive T2I agent system that is able to ask clarification questions and present its belief graph for the user to edit.
    • Novel human-agent interaction interface: We designed a new interface to best enable the clarification and belief graph interaction modalities. We have not seen these features in any T2I, or other generative media apps that are publicly live to date, signifying to us total uniqueness. Our human studies showed that at least 85% of raters found each component of the interface useful for their workflow, for us proving that these are both novel and useful.
    • Novel design of different T2I agents that enable the proposed interaction modalities. Please see §4 for the full details of the design principles and construction of those T2I agent prototypes (Ag1, Ag2, Ag3).
  2. 2.
    Our belief graphs significantly differs from classic belief states in the following ways:
    • Hardcoded predicates v.s. Automatically-generated predicates: Traditionally, constructing classic symbolic belief states requires a pre-defined set of predicates such as “on(a, b)”, “is_red(a)”, “at_position(robot, x, y, z)” and it is non-trivial to learn new predicates that can be used and generalized to new tasks (Pasula et al., 2007; Xia et al., 2019). Typically the pre-defined set of predicates are written by system developers and hardcoded into classic AI systems (Fikes and Nilsson, 1971). Our belief graphs do not require any pre-defined predicates. Instead, we propose to construct symbolic beliefs using a sequential in-context learning (ICL) method with LLMs (Algorithm 1). Our method can be generalized across a wide range of T2I tasks and achieve high performance (see our comprehensive results on Coco, Imageinwords, DesignBench).
    • Applications: To the best of our knowledge, classic symbolic beliefs are mostly used for robot planning, and we are the first to use symbolic belief graphs to assist T2I tasks.
    • Data structure: Because of the application to planning, a symbolic world state is usually implemented and stored as a set or list of literals, i.e., atoms or negation of atoms where atoms are instantiated predicates (Alkhazraji et al., 2020; Garrett et al., 2020a, b), so that whenever an action is applied, the agent can apply transition by adding and deleting items in the set according to the precondition and effect of the action.
      For T2I tasks, it is more convenient to use a graph to represent the world state associated with an image, since entities and relations naturally form a set of nodes (entities) and edges (relations between entities). Each component of the graph can also have probabilities, making it easy to turn a world state into a belief state using the same data structure. Hence we represent T2I agent beliefs using graphs. The agent can directly update the graph for each transition instead of going through a set or list.
    • Interpretability and controllability: The graph structure makes our agent belief more interpretable than traditional belief states, since we can visualize and progressively disclose the graph to the human user. Moreover, each node or relation in the belief graph has associated descriptions, making it easy for the user to understand and potentially edit every component of the belief graph. In our human studies, about 85% of raters found the belief graph useful. To the best of our knowledge, our work is the first to use the graph-based belief state for human-AI interaction.
  3. 3.
    Automated evaluation of T2I agents: We propose a novel automated evaluation approach for T2I agents using self-play. The agent interacts with a simulated user that has access to the original image and its long caption. See §5.1 and §E.4 for the full details of how the simulated user is constructed. This evaluation pipeline is easy to use and can help the future development of T2I agents.
  4. 4.
    DesignBench: We envision that a significant fraction of T2I users are artists and designers, and it is important to ensure that T2I agents are evaluated for these use cases. Hence we create DesignBench, featuring photo-realism, animation and multiple styles with short and long captions. DesignBench can be directly plugged into our automated evaluation to streamline the evaluation process.
Appendix C Formalism of the agent and its objective
We define an interactive T2I agent as a $\langle B,A,O,\tau,\pi\rangle$ tuple, where we have
  • $S$: a representation space of images,
  • $B$: a space of agent beliefs,
@@ Appendix C Formalism of the agent and its objective
For each user-initiated interaction, we assume that there exists a specific intent $s\in S$, where $S$ is the space of all possible user intents. For a T2I task, we assume that the intent is the image the user would like to generate, and the intent stays the same throughout the interaction with an agent. We discuss more about the validity of this assumption in §6.
@@
Appendix D Visualization of Multi-Turn Agent-User Dialogs and Generated Images
@@
Figure 7: Agent Generated Image Outputs on DesignBench: a chart of the generated image outputs of the four main Agent types in comparison to the goal image. Each column displays the output of a different agent and the right most column shows the goal image that the agents aimed to recreate. Each agent was provided with the same starting prompt and iterated for 15 turns, with the exception of the "T2I" agent column which produces an image from the starting prompt. Ag1, Ag2 and Ag3 refer to the Agents described in §E. Each agent uses the same T2I model to produce the final image. The goal images displayed here are from our DesignBench dataset described in the experiments section.
Figure 8: Agent Generated Image Outputs on DesignBench (Continued): a chart of the generated image outputs of the four main Agent types in comparison to the goal image. Each column displays the output of a different agent and the right most column shows the goal image that the agents aimed to recreate. Each agent was provided with the same starting prompt and iterated for 15 turns, with the exception of the "T2I" agent column which produces an image from the starting prompt. Ag1, Ag2 and Ag3 refer to the Agents described in §E. Each agent uses the same T2I model to produce the final image. The goal images displayed here are from the DesignBench dataset described in the experiments section.
Figure 9: Agent Generated Image Outputs (Coco-Captions Validation): a chart of the generated image outputs of the four main Agent types in comparison to the goal image. Each column displays the output of a different agent and the right most column shows the goal image that the agents aimed to recreate. Each agent was provided with the same starting prompt and iterated for 15 turns, with the exception of the "T2I" agent column which produces an image from the starting prompt. Ag1, Ag2 and Ag3 refer to the Agents described in §E. Each agent uses the same T2I model to produce the final image. The goal images displayed here are from the Coco-Captions Chen et al. (2015) dataset described in the experiments section.
@@
Appendix E Implementation details of agent prototypes
@@
  • Ag3: Principle-prompted Agent. This agent generates questions directly from the conversation history based on the principles introduced in §4.1.1. The question asking strategy of Ag3 relies solely on the implicit knowledge and reasoning capabilities of the underlying LLM.
E.1 Implementation of agent beliefs
@@
Prompt Merging. An LLM (§E.11) summarizes the latest interaction, encapsulating the agent’s question and the user’s response into a concise textual representation. This step distills the essential information exchanged during the interaction. Another LLM (§E.12) merges the existing merged prompt (containing the accumulated information from previous interactions) and the summarized interaction at the current turn. This creates an updated prompt that reflects the evolving understanding of the user’s intent.
Belief Parsing. See an example of the belief graph fig. 11. We employ three specialized parsers trained via in-context learning (ICL): entity parser (§E.8) analyzes the user prompt to identify and extract a list of relevant entities.; attribute parser (§E.9) takes user prompt and an entity as the input to extract a list of attributes associated with that entity; relation parser (§E.10) takes the user prompt and a list of entities as input and identifies relationships between those entities. Each entity is associated with meta information like name, importance to ask score, description, probability of appearing, a list of attributes like color, position, etc  [^10] Name is a unique identifier for the entity; Importance to ask score: A numerical value indicating the entity’s perceived importance in satisfying the user’s request. Entities with higher scores are prioritized during question generation, as they are likely to reduce uncertainty and contribute significantly to the final image; Description provides a textual description of the entity; probability of appearing estimates likelihood of the entity being present in the generated image; Attributes is for understanding the detailed attributes of the entities.. Each attribute contains meta information like name, importance to ask score, a list of possible values for the attribute along with their associated probabilities, etc. Each relation includes meta information such as: name, description, spatial relation, importance to ask score, entity 1 and entity 2, whether the relation is bidirectional, etc.
@@ E.2 Implementation of action
    Here $IS,P,Ent$ represents importance to ask score, probability of appearing, and entropy of the probabilities respectively and $e,a,c,r$ represents entity, attribute, candidate list, relation respectively.
  • Ask Important Clarification Question based on belief ($AICQ_{B}$): This strategy leverages the structured information within the belief. We provide the LLM with the user prompt, conversation history, and the current belief, utilizing an ICL prompt (§E.14) to guide question generation. The LLM then formulates a clarification question aimed at eliciting information about key features of the image, naturally prioritizing those with higher Importance to ask score within the belief.
  • Ask Important Clarification Question directly ($AICQ_{base}$): This strategy relies on the LLM’s inherent ability to identify important aspects of the user prompt and conversation history. The LLM (§E.13) generates an important clarification question based on its implicit understanding of the user’s needs, without explicitly relying on the structured information in the belief.
Ag1 employs $MHIS$ strategy for question generation. This strategy leverages the importance scores assigned to entities, attributes, and relations within the belief graph. It identifies the element with the highest heuristic importance score and formulates a question aimed at eliciting further information about that specific element. The question is then verbalized using the LLM described in Section §E.15.
@@
E.3 Implementation of belief transition
E.4 User simulation
@@
  • Mirroring the $AICQ_{B}$ strategy, the simulator takes the ground truth prompt, conversation history, and its current belief graph as input. It then leverages an ICL prompt (see §E.14) to generate a response to the agent’s question. This ensures that the simulator’s answers are consistent with its internal belief graph and the ongoing conversation.
E.5 Ag1: Heuristic Score Agent
The Heuristic score agent leverages the importance scores and probabilities within the belief graph to guide its question-asking strategy. The underlying principle is to identify and inquire about the entity, attribute, or relation that exhibits both high importance and high uncertainty. This aligns with the uncertainty reduction principle discussed in §4.1.1, which emphasizes minimizing uncertainty through targeted questioning. To achieve this, we define a heuristic importance score as formulated in Equation 1, and the agent then selects the attribute or relation with the highest heuristic importance score as the focus of its inquiry. To facilitate easy answering, we utilize an LLM to generate user-friendly questions with multiple-choice options. For example, the agent might ask: What color of the rabbit do you have in mind? a. black , b. white, c. brown. d. unkown. If none of these options , what color of the rabbit do you have in mind?. This format allows users to simply select the most appropriate option or provide their own answer if needed.
Here’s a summary of Ag1’s implementation: (i) Belief Representation: The agent’s belief comprises the merged prompt and the current belief graph. (ii) Select Action: $MHIS$ strategy is employed to identify the attribute or relation of interest based on the heuristic importance score. (iii) Verbalize Action: An LLM (§E.15) is used to generate a clear and concise question about the selected attribute or relation. (iv) Answer Question: The user simulator provides an answer to the agent’s question, mimicking human response behavior. (v) Transition: The agent updates the merged prompt with the new information, re-generates the belief graph based on the updated prompt, and applies the post-processing logic outlined in §E.3 to ensure consistency and prevent redundancy.
E.6 Ag2: Belief-prompted Agent
The Ag2 agent incorporates the belief graph into its decision-making process but adopts a different approach compared to Ag1. Instead of relying on a heuristic score, Ag2 leverages the full capacity of an LLM to generate questions. It provides the LLM with comprehensive information, including the merged prompt, belief graph, and conversation history, allowing the LLM to formulate the most informative questions possible. To guide the LLM towards generating effective questions, we incorporate specific instructions in the prompt, emphasizing the following principles: The question should be as concise and direct as possible. The question should aim to obtain the most information about the style, entities, attributes, spatial layout and other contents of the image. Remember to ask for information that are critical to knowing the critical details of the image that is important to the user. The question should reduce your uncertainty about the user intent as much as possible.
Here’s a summary of Ag2’s implementation: (i) Belief Representation: The same as Ag1, the agent’s belief consists of the merged prompt and the current belief graph. (ii) Select Action: $AICQ_{B}$ strategy is employed, which leverages an LLM to generate a question based on the comprehensive input information. (iii) Verbalize Action: Since the LLM directly generates the question, no separate verbalization step is required. (iv) Answer Question: The user simulator provides an answer to the agent’s question. (v) Transition: The agent updates the merged prompt with the new information and re-generates the belief graph based on the updated prompt.
E.7 Ag3: Principle-prompted Agent
A simple and effective implementation of LLM-based multi-modal dialogue systems is to use the context to store the history of conversations between the system and the user, and directly generate the next response based on the context.
To align with the principles outlined in §4.1.1, we guide the LLM’s question generation with a prompt (§E.13) that emphasizes all principles: Based on the original prompt and chat history please provide a question to ask about the image. The question should be as concise and direct as possible. The question should aim to learn more about the attributes and contents of the image, the objects, the spatial layout, and the style. The prompt also includes the history of conversation. This strategy aims to generate questions that are easy for users to understand and answer, while effectively reducing the agent’s uncertainty about the desired image.
Here’s a summary of Ag3’s implementation: (i) Belief Representation: The same as Ag1 and Ag2, the agent’s belief consists of the merged prompt and the current belief graph. (ii) Select Action: $AICQ_{base}$ strategy is employed, which leverages an LLM to generate a question based on the conversation history. (iii) Verbalize Action: The LLM directly generates the question, so no separate verbalization step is needed. (iv) Answer Question: The user simulator provides an answer to the agent’s question. (v) Transition: The same as Ag2, the agent updates the merged prompt with the new information and re-generates the belief graph based on the updated prompt.
E.8 Entity Parser Prompt Instruction

### v2

@@ 7.4 Human Studies on the Agent Interface
Table 2: Perceived helpfulness of proposed features, shown as % of users out of 143 human raters.
@@ 8 Discussion and Conclusion
This work introduces a design for agents that assist users in generating images through an interactive process of proactive question asking and belief graph refinement. By dynamically updating its understanding of the user’s intent, the agent facilitates a collaborative approach to image generation with fewer user frustrations. Moreover, presenting the agent’s belief graph can be a generalizable method for AI transparency, which is an important factor given the increasing complexity of modern AI models.
Modular design. Our agent prototypes are highly modular: they use frozen T2I models to generate images based on prompts updated by the agents. This allows for seamless integration of improved off-the-shelf T2I models as they become available, boosting system performance without requiring further adaptation.
While T2I prompt-image alignment errors can limit the effectiveness of our proposed agents, our experiment in §E.2 demonstrates that by leveraging agent-user QA pairs to evaluate image fidelity, we can filter out misaligned images across N seeds, thereby mitigating the impact of these errors. Note that the agents’ T2T scores in Table 1 (92%+) ablate the T2I model, showing that they have already reached high alignment on the caption level, which is bound to increase with improved T2I capabilities.
Culturally competent models: There has been a growing use of generative AI models by a broader section of society, which is only expected to continue to grow. With the varied preferences and intents across diverse user bases, this work is an example of steerable models that can adapt better to diverse user needs. Moreover, different groups of people may perceive harmfulness of content differently, so learning more about the user through clarification questions can mitigate risks, make models safer, and pave the way to more inclusive generative AI systems.
Future work. It would be interesting to explore generating images directly from belief graphs and fine-tuning VLMs on image-text interleaved multi-turn trajectory data. This may require a) collecting data such as gold-standard trajectories or annotations on the quality of trajectories of human-agent conversations and b) approaches to fine-tune the model on multi-turn trajectories of images and text, which can potentially improve the performance of the agent.
Impact Statement
Our proposed T2I agents are equipped with better tools (belief graphs) for interpretability and controllability. Presenting the agent’s belief graph can be a generalizable method for AI transparency, which is an important factor given the increasing complexity of modern AI models.
By asking clarification questions, our proposed agents may enable a more customizable and personalized content creation experience. Because different groups of people may perceive harmfulness of contents differently, learning more about the user through clarification questions can potentially mitigate risks of generating contents that can be offensive to each specific user.
@@ Acknowledgements
We would like to thank Jason Baldridge and Zoubin Ghahramani for insightful discussions on multi-turn T2I and belief states, Mahima Pushkarna for the help and consultation on user study. We would also like to thank Richard Song, Noah Fiedel and anonymous reviewers for feedback on the paper.
@@ Appendix A Author contributions
Appendix B Formalism of the Agent and its Objective
We define an interactive agent as a $\langle B,A,O,\tau,\pi\rangle$ tuple, where we have
  • $S$: a state space where each state represents an intent,
  • $B$: a space of agent belief states,
@@ Appendix B Formalism of the Agent and its Objective
For each user-initiated interaction, we assume that there exists a specific intent $s\in S$ in hindsight, where $S$ is the space of all possible user intents. For a T2I task, this intent is the image the user would like to generate at the end of the interaction.
@@
Appendix C Visualization of Multi-Turn Agent-User Dialogs and Generated Images
@@
Figure 6: Agent Generated Image Outputs on DesignBench: a chart of the generated image outputs of the four main Agent types in comparison to the goal image. Each column displays the output of a different agent and the right most column shows the goal image that the agents aimed to recreate. Each agent was provided with the same starting prompt and iterated for 15 turns, with the exception of the "T2I" agent column which produces an image from the starting prompt. Ag1, Ag2 and Ag3 refer to the Agents described in §D. Each agent uses the same T2I model to produce the final image. The goal images displayed here are from our DesignBench dataset described in the experiments section.
Figure 7: Agent Generated Image Outputs on DesignBench (Continued): a chart of the generated image outputs of the four main Agent types in comparison to the goal image. Each column displays the output of a different agent and the right most column shows the goal image that the agents aimed to recreate. Each agent was provided with the same starting prompt and iterated for 15 turns, with the exception of the "T2I" agent column which produces an image from the starting prompt. Ag1, Ag2 and Ag3 refer to the Agents described in §D. Each agent uses the same T2I model to produce the final image. The goal images displayed here are from the DesignBench dataset described in the experiments section.
Figure 8: Agent Generated Image Outputs (Coco-Captions Validation): a chart of the generated image outputs of the four main Agent types in comparison to the goal image. Each column displays the output of a different agent and the right most column shows the goal image that the agents aimed to recreate. Each agent was provided with the same starting prompt and iterated for 15 turns, with the exception of the "T2I" agent column which produces an image from the starting prompt. Ag1, Ag2 and Ag3 refer to the Agents described in §D. Each agent uses the same T2I model to produce the final image. The goal images displayed here are from the Coco-Captions Chen et al. (2015) dataset described in the experiments section.
@@
Appendix D Implementation Details of Agent Prototypes
@@
  • Ag3: Principle-prompted Agent. This agent generates questions directly from the conversation history based on the principles introduced in §5.2. The question asking strategy of Ag3 relies solely on the implicit knowledge and reasoning capabilities of the underlying LLM.
A summary of each agent is located in §6.
D.1 Implementation of Agent Beliefs
@@
Prompt Merging. An LLM (§D.8) summarizes the latest interaction, encapsulating the agent’s question and the user’s response into a concise textual representation. This step distills the essential information exchanged during the interaction. Another LLM (§D.9) merges the existing merged prompt (containing the accumulated information from previous interactions) and the summarized interaction at the current turn. This creates an updated prompt that reflects the evolving understanding of the user’s intent.
Belief Parsing. See an example of the belief graph in fig. 10. We employ three specialized parsers trained via in-context learning (ICL): entity parser (§D.5) analyzes the user prompt to identify and extract a list of relevant entities.; attribute parser (§D.6) takes user prompt and an entity as the input to extract a list of attributes associated with that entity; relation parser (§D.7) takes the user prompt and a list of entities as input and identifies relationships between those entities. Each entity is associated with meta information like name, importance to ask score, description, probability of appearing, a list of attributes like color, position, etc  [^6] Name is a unique identifier for the entity; Importance to ask score: A numerical value indicating the entity’s perceived importance in satisfying the user’s request. Entities with higher scores are prioritized during question generation, as they are likely to reduce uncertainty and contribute significantly to the final image; Description provides a textual description of the entity; probability of appearing estimates likelihood of the entity being present in the generated image; Attributes are for understanding the detailed attributes of the entities.. Each attribute contains meta information like name, importance to ask score, a list of possible values for the attribute along with their associated probabilities, etc. Each relation includes meta information such as: name, description, spatial relation, importance to ask score, entity 1 and entity 2, whether the relation is bidirectional, etc.
@@ D.2 Implementation of Action Selection
    Here $IS,P,Ent$ represent importance to ask score, probability of appearing, and entropy of the probabilities respectively and $e,a,c,r$ represent entity, attribute, candidate list, relation respectively.
  • Ask Important Clarification Question based on belief ($AICQ_{B}$): This strategy leverages the structured information within the belief. We provide the LLM with the user prompt, conversation history, and the current belief, utilizing an ICL prompt (§D.11) to guide question generation. The LLM then formulates a clarification question aimed at eliciting information about key features of the image, naturally prioritizing those with higher Importance to ask score within the belief.
  • Ask Important Clarification Question directly ($AICQ_{base}$): This strategy relies on the LLM’s inherent ability to identify important aspects of the user prompt and conversation history. The LLM (§D.10) generates an important clarification question based on its implicit understanding of the user’s needs, without explicitly relying on the structured information in the belief.
Ag1 employs $MHIS$ strategy for question generation. This strategy leverages the importance scores assigned to entities, attributes, and relations within the belief graph. It identifies the element with the highest heuristic importance score and formulates a question aimed at eliciting further information about that specific element. The question is then verbalized using the LLM described in Section §D.12.
@@
D.3 User Simulation
@@
  • The simulator takes the ground truth prompt, conversation history, and its current belief graph as input. It then leverages an ICL prompt (see §D.11) to generate a response to the agent’s question. This ensures that the simulator’s answers are consistent with its internal belief graph and the ongoing conversation.
D.4 Implementation of Belief Transition
D.5 Entity Parser Prompt Instruction